OpenStudy (anonymous):

find fist and second derivate x^2 - 2x + 1/ x+1

OpenStudy (anonymous):

qoutient rule?

OpenStudy (anonymous):

So if marginal profit is P(x) = 3x^2 - 10x +7, and P'(x) =, 6x - 10, what do I do?

OpenStudy (vijay):

\[y' = 2(x-1)/(x+1) -(x-1)/ (x+1)^2\]

OpenStudy (vijay):

\[1st derivative (y') =(x^2 +2x - 3) / (x+1)^2 \]

OpenStudy (anonymous):

im still confused i undertand until y' but the other part where does it come from??

OpenStudy (vijay):

u mean the 2nd answer, if u factorise (x-1)/(x+1) from 1st answer , u will get

OpenStudy (anonymous):

oh i see it know so basically to get now the 2nd derivate use that and go from there

OpenStudy (vijay):

yes

OpenStudy (anonymous):

thanks